OpenStudy (anonymous):

30 = -4.4 square root x + 38

OpenStudy (anonymous):

\[30= -4.4\sqrt{x} + 38\]

OpenStudy (anonymous):

sqrt x = 2/1.1 x = 4 / 1.21

OpenStudy (anonymous):

did u got it or not

OpenStudy (anonymous):

can you explain how you got that? :/